OpenStudy (anonymous):

What is the equation of the line passing through (0, 2) and perpendicular to the line y = 4x – 3 in slope-intercept form?

OpenStudy (zzr0ck3r):

did u understand the last one?

OpenStudy (anonymous):

sort of.

OpenStudy (zzr0ck3r):

perp means m = -1/2 so same as before y = -1/2(x) + b what is b? use the point 2 = -1/2(0)+b b = 2 so y = (-1/2)x+2